# Homogeneous $\mathrm{G}_{2}$ and Sasakian instantons on the Stiefel $7$-manifold

### v1

@@ 4.1 Homogeneous ${\rm SO}(3)$-principal bundles
Consider $H=K={\rm SO}(3)$ and $G={\rm SO}(5)$. Since ${\rm SO}(3)$ is simple, up to conjugation there are two smooth homomorphisms $\phi_{i}\colon{\rm SO}(3)\to{\rm SO}(3),$ $i=0,1$. Namely, $\phi_{0}$ the trivial homomorphism and $\phi_{1}=1_{{\rm SO}(3)}$ the identity, according to Lemma 4.2, there exist precisely two [^3] Up to isomorphism of homogeneous bundles homogeneous ${\rm SO}(3)$-principal bundles $\pi_{i}\colon P_{i}=G\times_{\phi_{i}}{\rm SO}(3)\to V^{5,2}$, $i=0,1$.
@@
$h_{0}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},h_{1}=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},h_{2}=\begin{pmatrix}0&0&1\\ 0&-1&0\\ 1&0&0\end{pmatrix},\quad\text{and}\quad h_{3}=\begin{pmatrix}0&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix}$ (4.2)
@@
follows that $a_{i}=b_{i}=c_{i}=0$ for $i\in\{2,4,5,7\}$. Now, by considering $h_{2}$ and $h_{3},$ we obtain for $i=3,6$
@@
The curvature of $\alpha$ is $F_{\alpha}=y_{2}^{-1}(X^{25}+X^{36}+X^{47})\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10})$. The second affirmation in the Lemma follows from (2.6) since
@@
Remark 4.5 (Homogeneous ${\rm U}(1)$-connections).
@@
From item $(i)$, since $\alpha|_{\mathfrak{so}(3)}=0$, $\alpha$ has to be of the form $\alpha=a_{i}e^{i}\otimes e_{1}$, $i=1,\cdots,7.$ On the other hand, considering item $(ii)$ for ${\textsl{g}}_{0},{\textsl{g}}_{1}$ given in (4.2), we use the same argument used in the proof of Lemma 4.4 to obtain $\alpha(e_{i})=a_{i}e_{1}$ and
$\alpha(Ad({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1}))$ $=-a_{i}e_{1}$
$\alpha(Ad({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1}))$ $=a_{i}e_{1}.$
for $i=2,\cdots,7$. Therefore, $\alpha$ is characterised by
@@
Note that $F_{\beta}\in\Omega^{2}_{1}=\omega\otimes\Omega^{0}(M)$ (4.13).
@@
and the class is given by $[(u,s)]=\{(u\cdot r,r^{-1}s)\;|\;r\in{\rm SO}(3)\}.$ There exist a left ${\rm SO}(5)$-action on $P_{1}$ such that for each $v\in{\rm SO}(5)$ the bundle map $L_{v}\colon P_{1}\to P_{1}$ is $\pi_{1}$-equivariant i.e.
@@
According to (2.2), we fix the basis $\{e_{8},e_{9},e_{10}\}$ for $\mathfrak{so}(3)\subset\mathfrak{so}(5)$ such that $[e_{i},e_{j}]=\varepsilon_{ijk}e_{k}$, where $\varepsilon_{ijk}$ is the sign of the permutation $(i\;j\;k)$.
@@
Consider $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$ in the basis (2.2) given by $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$. From Lemma 4.3 (i), we have that $\alpha=\alpha_{0}+e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10},$ where $\alpha_{0}=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$, $i,j,k\in\{1,\dots,7\}$, we are going to check that $\alpha_{0}=0$. From Lemma 4.3 (ii), for any ${\textsl{g}}\in{\rm SO}(3)$ and $X\in\mathfrak{so}(5)$, we have ${\textsl{g}}\alpha(X){\textsl{g}}^{-1}=\alpha({\textsl{g}}X{\textsl{g}}^{-1})$. Taking $X=e_{i}$ for each $i\in\{1,\dots,7\}$, we denote
@@
These imply $-a_{2}=a_{3},-c_{2}=b_{3},c_{3}=b_{2}$. Analogously, for $h_{1}$, we have
@@
consequently, $-a_{2}=a_{3},c_{2}=b_{3},-b_{2}=c_{3}$, hence $c_{2}=b_{3}=c_{3}=b_{2}=0$. Similarly:
@@ 4.2 Homogeneous $\mathrm{G}_{2}$ and self dual contact instantons (SDCI)
For any Sasakian manifold, the bundle of $2$–forms decomposes into [portilla2023, Section 2.1]
@@
Consider $V^{5,2}$ endowed with the Sasakian structure given in Corollary 2.8. There exist an unique invariant ASDCI on the trivial homogeneous ${\rm SO}(3)$–principal bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4).
@@
Consider the connection (4.8) on $V^{5,2}$, with respect to the $\mathrm{G}_{2}$–structure given in (2.11). Then we gave
  • The ${\rm SO}(3)$–connection in (4.8) on the homogeneous principal bundle $P_{1}\to V^{5,2}$ [cf. Lemma 4.6] satisfies $F_{\alpha}\wedge\psi=0$ for any $y_{1}>0$, i.e., there exist an unique ${\rm SO}(5)$–invariant connection which is simultaneously $\mathrm{G}_{2}$ and SDCI, given by (4.11).
@@
4.3 Stability of SDCI and $\mathrm{G}_{2}$-instantons
Now, we analyse the Yang-Mills condition for instanton (4.8), for a detailed study on stability of Yang-Mills connections we refer to [bourguignon1981stability]. On the space of connections $\mathcal{A}$, the global inner product $(\alpha,\beta)=\int_{M}\langle\alpha\wedge\ast\beta\rangle$ allow us define the Yang-Mills functional as the $L^{2}$-norm of the curvature:
@@ 4.3 Stability of SDCI and $\mathrm{G}_{2}$-instantons
A critical point of (4.14) is called a Yang-Mills (YM) connection, i.e. $\nabla$ satisfies the so-called Yang-Mills equation $d^{\ast}_{\nabla}F_{\nabla}=0$. Taking the exterior derivative $d_{\nabla}$ in (1.2) and using the Bianchi identity, one note that a solution of the instanton equation (1.2) satisfy the YM-equation with torsion:
@@
Note that neither $\sigma=\eta\wedge d\eta$ nor $\sigma=\varphi$ are closed, consequently, the torsion term in the above equation is no automatically zero, however, we have that SDCI and ASDCI are Yang-Mills connections [luis2024Weitzenbock, Section 5]. With respect to the $\mathrm{G}_{2}$-structure (2.23), we have the following:
@@
Let $\alpha$ be the unique $\mathrm{G}_{2}$-instanton (4.8) on the homogeneous ${\rm SO}(3)$-principal bundle $P_{1}$. Then, $\alpha$ is a critical point for the Yang-Mills energy for any coclosed $\mathrm{G}_{2}$-structure (2.23).
@@
For the Sasakian case, we can prove that SDCI are local minimizers of the YM-functional (4.14). In order to prove this, note that the contact structure induces an orthogonal decomposition (cf. [portilla2023, Lemma 2.3]) of the curvature $2$-form ${\rm F}_{\nabla}$ into
${\rm F}_{H}=i_{\xi}(\eta\wedge{\rm F}_{\nabla})\quad\text{and}\quad{\rm F}_{V}=\eta\wedge i_{\xi}{\rm F}_{\nabla},$
where $F_{H}$ is the horizontal part and $F_{V}$ the vertical part of $F_{\nabla}$. Consequently, we have $\|{\rm F}_{\nabla}\|^{2}=\|{\rm F}_{H}\|^{2}+\|{\rm F}_{V}\|^{2}$ and denote
${\rm F}_{H}^{+}=\frac{1}{2}({\rm F}_{H}+\ast(\sigma\wedge{\rm F}_{H}))\quad\text{and}\quad{\rm F}_{H}^{-}=\frac{1}{2}({\rm F}_{H}-\ast(\sigma\wedge{\rm F}_{H})),$ (4.15)
According to the splitting (3.2), we have ${\rm F}_{H}={\rm F}_{8}+{\rm F}_{6}+{\rm F}_{1}\in\Omega_{8}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{6}^{2}({\mathfrak{g}}_{E})\oplus\Omega_{1}^{2}({\mathfrak{g}}_{E})$, then we observe that ${\rm F}_{H}^{+}={\rm F}_{8}-\frac{1}{2}{\rm F}_{1}$ and ${\rm F}_{H}^{-}={\rm F}_{6}+\frac{3}{2}{\rm F}_{1}$, hence
$\|{\rm F}_{H}\|^{2}$ $=\langle{\rm F}_{H}^{+}+{\rm F}_{H}^{-},{\rm F}_{H}^{+}+{\rm F}_{H}^{-}\rangle=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+2\langle{\rm F}_{H}^{+},{\rm F}_{H}^{-}\rangle$
$=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2},$
consequently, the YM-functional (4.14) splits as
$\mathcal{YM}(\nabla)=\|{\rm F}_{H}^{+}\|^{2}+\|{\rm F}_{H}^{-}\|^{2}+\|{\rm F}_{V}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}.$ (4.16)
In the Sasakian context, the charge of the connection $\nabla$ is defined by
$\kappa(\nabla)=\displaystyle\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}^{2})\wedge\eta\wedge\omega.$ (4.17)
Notice that the charge (4.17) is determined by the horizontal part of $F_{\nabla}$ [portilla2023, Proposition 3.13]:
$\kappa(\nabla)=\|{\rm F}_{H}^{+}\|^{2}-\|{\rm F}_{H}^{-}\|^{2}.$ (4.18)
Using (4.16) and (4.18), we have that the YM-mills action is given by
$\mathcal{YM}(\nabla)$ $=\kappa(\nabla)+2\|{\rm F}_{H}^{-}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2}$
$=-\kappa(\nabla)+2\|{\rm F}_{H}^{+}\|^{2}-\frac{3}{2}\|{\rm F}_{1}\|^{2}+\|{\rm F}_{V}\|^{2},$
hence, we have
$\mathcal{YM}(\nabla)=\begin{cases}\kappa(\nabla),&\nabla\in\Omega^{2}_{8}({\mathfrak{g}}_{E})\\ -\kappa(\nabla),&\nabla\in\Omega^{2}_{6}({\mathfrak{g}}_{E})\\ 2\kappa(\nabla),&\nabla\in\Omega^{2}_{1}({\mathfrak{g}}_{E}).\end{cases}$ (4.19)
We recall that if $M$ is a contact manifold, a $k$-form $\alpha\in\Omega^{k}(M)$ is called transversal form if $i_{\xi}\alpha=0$, we denote the space of transversal forms by $\Omega^{k}_{H}(M)$. Consequently, the space of $k$-forms splits as
$\Omega^{k}(M)=\Omega^{k}_{H}(M)\oplus\eta\wedge\Omega^{k-1}_{H}(M).$ (4.20)
More generally, consider a Riemannian manifold $M$ endowed with a Riemannian foliation $\mathcal{F}\subset M$ and consider a principal $G$-bundle $\pi\colon P\to M$. $P$ is said to be a foliated principal bundle if there exists a $G$-invariant foliation $\widetilde{\mathcal{F}}$ on $P$ such that the tangent distribution $T\widetilde{\mathcal{F}}\subset TP$ projects isomorphically onto $T\mathcal{F}$ under $\pi_{\ast}\colon TP\to TM$. Accordingly, a vector bundle is called foliated if its associated principal bundle is foliated. A local section $s\colon U\subset M\to P$ is said basic section if $s_{\ast}\colon TU\to TP$ maps $T\mathcal{F}$ to $T\widetilde{\mathcal{F}}$. It is known that every foliated principal bundle admits local basic sections [molino1988riemannian, Proposition 2.7]. We denote the space of basic sections by $\Gamma(E)_{B}\subset\Gamma(E)$. In the context of contact manifolds, $\mathcal{F}$ is naturally given by the Reeb foliation, if $\nabla\in\Gamma(E)$ is a connection with $i_{\xi}F_{\nabla}=0$, then the horizontal lift of $\mathcal{F}$ with respect to $\nabla$ is an integrable distribution and gives $E$ a foliated structure for which $\nabla$ is basic connection [molino1988riemannian], consequently, we have a natural split of the bundle of $E$-valued $k$-forms analogous to (4.20).
On the underlying bundle $E$, considering only transverse deformations, i.e., $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}_{H}({\mathfrak{g}}_{E})$, in [portilla2023, Lemma 3.14] is shown that the charge vary by
$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}.$ (4.21)
The following result emulates [portilla2023, Lemma 3.14], the proof proceeds with the same reasoning, but we emphasise that, first, this result is not exclusive to the Stiefel manifold, but rather applies to any Sasakian manifold. And secondly, in [portilla2023] the result is announced for transverse deformations of the Chern connection [portilla2023, Corollary 3.15], but actually this result holds in a more general context as follow:

### v2

@@ 4.1 Homogeneous ${\rm SO}(3)$-principal bundles
Consider $H=K={\rm SO}(3)$ and $G={\rm SO}(5)$. Since ${\rm SO}(3)$ is simple, up to conjugation there are two smooth homomorphisms $\phi_{i}\colon{\rm SO}(3)\to{\rm SO}(3),$ $i=0,1$. Namely, $\phi_{0}$ the trivial homomorphism and $\phi_{1}=1_{{\rm SO}(3)}$ the identity homomorphism, according to Lemma 4.2, there exist precisely two [^3] Up to isomorphism of homogeneous principal bundles homogeneous ${\rm SO}(3)$-principal bundles $\pi_{i}\colon P_{i}=G\times_{\phi_{i}}{\rm SO}(3)\to V^{5,2}$, $i=0,1$.
@@
$h_{0}=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},h_{1}=\begin{pmatrix}0&1&0\\ -1&0&0\\ 0&0&1\end{pmatrix},h_{2}=\begin{pmatrix}0&0&1\\ 0&-1&0\\ 1&0&0\end{pmatrix}\quad\text{and}\quad h_{3}=\begin{pmatrix}0&0&1\\ 0&1&0\\ -1&0&0\end{pmatrix}$ (4.2)
@@
it follows that $a_{i}=b_{i}=c_{i}=0$ for $i\in\{2,4,5,7\}$. Now, by considering $h_{2}$ and $h_{3},$ we obtain for $i=3,6$
@@
The curvature of $\alpha$ is $F_{\alpha}=y_{2}^{-1}(X^{25}+X^{36}+X^{47})\otimes(b_{8}e_{8}+b_{9}e_{9}+b_{10}e_{10})$. The second statement in the Lemma follows from (2.6) since
@@
Lemma 4.5 (Homogeneous ${\rm U}(1)$-connections).
There is a unique invariant ${\rm U}(1)$-connections on $P=V^{5,2}\times{\rm U}(1)$, given by
$\beta=a_{1}e^{1}\otimes e_{1}$ (4.5)
Furthermore, the curvature $F_{\alpha}$ of the connection $\alpha$ belongs to $\Omega^{2}_{1}$ cf. (3.4).
Proof.
@@
From item $(i)$, since $\alpha|_{\mathfrak{so}(3)}=0$, $\alpha$ has to be of the form $\alpha=a_{i}e^{i}\otimes e_{1}$, $i=1,\cdots,7.$ On the other hand, considering item $(ii)$ for ${\textsl{g}}_{0},{\textsl{g}}_{1}$ given in (4.2), we use the same argument used in the proof of Lemma 4.4 to obtain $\alpha(e_{i})=a_{i}e_{1}$ and for $i=2,\cdots,7$
$\alpha(\mathrm{Ad}({\textsl{g}}_{0})(e_{i}))=\alpha({\textsl{g}}_{0}e_{i}{\textsl{g}}_{0}^{-1}))$ $=-a_{i}e_{1}$
$\alpha(\mathrm{Ad}({\textsl{g}}_{1})(e_{i}))=\alpha({\textsl{g}}_{1}e_{i}{\textsl{g}}_{1}^{-1}))$ $=a_{i}e_{1}.$
Therefore, $\alpha$ is characterised by
@@
Note that $F_{\beta}\in\Omega^{2}_{1}=\omega\otimes\Omega^{0}(M)$ (4.14). ∎
@@
and the class is given by $(u,s)]=\{(u\cdot r,r^{-1}s)\;|\;r\in{\rm SO}(3)\}.$ There exist a left ${\rm SO}(5)$-action on $P_{1}$ such that for each $v\in{\rm SO}(5)$ the bundle map $L_{v}\colon P_{1}\to P_{1}$ is $\pi_{1}$-equivariant, i.e.,
@@
According to (2.2), we fix the basis $\{e_{8},e_{9},e_{10}\}$ for $\mathfrak{so}(3)\subset\mathfrak{so}(5)$ such that $e_{i},e_{j}]=\varepsilon_{ijk}e_{k}$, where $\varepsilon_{ijk}$ is the sign of the permutation (ijk)(i\;j\;k).
@@
Consider $\alpha:\mathfrak{so}(5)\rightarrow\mathfrak{so}(3)$ in the basis (2.2) given by $\alpha=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$. From Lemma 4.3 (i), we have that $\alpha=\alpha_{0}+e^{8}\otimes e_{8}+e^{9}\otimes e_{9}+e^{10}\otimes e_{10},$ where $\alpha_{0}=a_{i}e^{i}\otimes e_{8}+b_{j}e^{j}\otimes e_{9}+c_{k}e^{k}\otimes e_{10}$, with $i,j,k\in\{1,\dots,7\}$, we are going to conclude that $\alpha_{0}=0$. From Lemma 4.3 (ii), for any ${\textsl{g}}\in{\rm SO}(3)$ and $X\in\mathfrak{so}(5)$, we have ${\textsl{g}}\alpha(X){\textsl{g}}^{-1}=\alpha({\textsl{g}}X{\textsl{g}}^{-1})$. Taking $X=e_{i}$ for each $i\in\{1,\dots,7\}$, we denote
@@
These imply $a_{2}=a_{3},-c_{2}=b_{3},c_{3}=b_{2}$. Analogously, for $h_{1}$, we have
@@
consequently, $a_{2}=a_{3},c_{2}=b_{3},-b_{2}=c_{3}$, hence $c_{2}=b_{3}=c_{3}=b_{2}=0$. Similarly:
@@ 4.2 Homogeneous $\mathrm{G}_{2}$ and self dual contact instantons (SDCI)
For any Sasakian manifold, the bundle of 22–forms decomposes into [portilla2023, Section 2.1]
@@
Consider $V^{5,2}$ endowed with the Sasakian structure given in Corollary 2.8. There exist a unique invariant ASDCI on the trivial homogeneous ${\rm SO}(3)$–principal bundle $P_{0}\to V^{5,2}$ (cf. Lemma 4.4).
@@
Consider the connection (4.9) on $V^{5,2}$, with respect to the $\mathrm{G}_{2}$–structure given in Corollary (2.8). Then, we have
  • The ${\rm SO}(3)$–connection in (4.9) on the homogeneous principal bundle $P_{1}\to V^{5,2}$ [cf. Lemma 4.6] satisfies $F_{\alpha}\wedge\psi=0$ for any $y_{1}>0$, i.e., there exist a unique ${\rm SO}(5)$–invariant connection which is simultaneously $\mathrm{G}_{2}$ and SDCI, given by (4.12).
@@
4.3 Yang–Mills connections
Now, we analyse the Yang–Mills condition for instanton (4.9), for a detailed study on stability of Yang–Mills connections we refer to [bourguignon1981stability]. On the space of connections $\mathcal{A}$, the global inner product $(\alpha,\beta)=\int_{M}\langle\alpha\wedge\ast\beta\rangle$ allow us to define the Yang–Mills functional as the $L^{2}$-norm of the curvature:
@@ 4.3 Yang–Mills connections
A critical point of (4.15) is called a Yang–Mills (YM) connection, i.e. $\nabla$ satisfies the so-called Yang–Mills equation $d^{\ast}_{\nabla}F_{\nabla}=0$. Taking the exterior derivative $d_{\nabla}$ in (1.2) and using the Bianchi identity, one notes that a solution of the instanton equation (1.2) satisfy the YM-equation with torsion:
@@
Note that neither $\sigma=\eta\wedge d\eta$ nor $\sigma=\varphi$ are closed, consequently, the torsion term in the above equation is not automatically zero, however, we have that SDCI and ASDCI are Yang–Mills connections [luis2024Weitzenbock, Section 5]. With respect to the $\mathrm{G}_{2}$-structure (2.23), we have the following:
@@
Let $\alpha$ be the unique $\mathrm{G}_{2}$-instanton (4.9) on the homogeneous ${\rm SO}(3)$-principal bundle $P_{1}$. Then, $\alpha$ is a critical point for the Yang–Mills energy for any coclosed $\mathrm{G}_{2}$-structure (2.23).
@@
5 Infinitesimal deformations and rigidity of $\mathrm{G}_{2}$-instantons
In this section, we extend the spinorial characterisation of infinitesimal deformations of $\mathrm{G}_{2}$–instantons obtained in [singhal2022] to the setting of coclosed $\mathrm{G}_{2}$–structures. As an application, we show that for certain values of the parameter $y_{2}$ appearing in (4.12), the corresponding $\mathrm{G}_{2}$–instanton is rigid. Throughout this section, $(M^{7},\varphi)$ denotes a 77–dimensional manifold endowed with a coclosed $\mathrm{G}_{2}$–structure, and $\zeta$ the unit spinor associated with $\varphi$.
Let us recall, consider $\zeta$ the unit spinor associated to $\varphi$ and the induced spinor representation $\Delta_{7}\simeq\langle\zeta\rangle\oplus TM$, the Clifford product of a vector field $Y$ and a spinor $(f,Z)$ is (see [karigiannis2006some])
$Y\cdot(f,Z)=(-\langle Y,Z\rangle,fY+Y\times Z),$ (5.1)
where $(f,0)=f\zeta$ and $Y\times Z=(\varphi(Y,Z,\cdot))^{\sharp}$. The Clifford product (5.1) is extended to a $k$-form $\varepsilon$ by
$\varepsilon\cdot(f,Y)=\frac{1}{k!}\varepsilon_{i_{1},\dots i_{k}}e_{i_{1}}\cdot(e_{i_{2}}\cdot(\dots\cdot(e_{i_{k}}\cdot(f,Z)\dots))$ (5.2)
Hence, we have the characterisation of the instanton condition (1.3) in the $\mathrm{G}_{2}$ case in terms of spinors:
Definition 5.1.
Let $P\to M$ be a principal bundle. A connection $A$ on a principal bundle $P$ is called a $\mathrm{G}_{2}$–instanton if its curvature satisfies $F_{A}\cdot\zeta=0,$ where $\cdot$ denotes Clifford multiplication (5.2).
5.1 Infinitesimal deformation of $\mathrm{G}_{2}$-instantons
Definition 5.2.
Let $A$ be a $\mathrm{G}_{2}$-instanton on a principal bundle $P\to M$. An infinitesimal deformation of $A$ is a one–form $\varepsilon\in\Omega^{1}(M,\mathrm{Ad}(P))$ satisfying
$(d_{A}\varepsilon)\cdot\zeta=0,\qquad d_{A}^{\ast}\varepsilon=0.$ (5.3)
Let $\nabla^{g}$ be the Levi–Civita connection of $g$ associated to $\varphi$, we consider the one-parameter family of connections
$\nabla^{t}=\nabla^{g}+\frac{t}{2}\,g^{-1}H_{\varphi}\quad\text{for}\quad t\in\mathbb{R},$ (5.4)
and
$H_{\varphi}=\frac{\tau_{0}}{6}\,\varphi-\tau_{3}.$
For $t=1$, it turns out that $\nabla^{1}$ is the unique connection with parallel skew-symmetric torsion $H_{\varphi}$ that makes $\varphi$ parallel [friedrich2002]. The connection $\nabla^{t}$ induces a connection on the spinor bundle $\not${S} such that:
[…]
Definition 5.3.
With respect to the Clifford multiplication $\cdot\colon TM\times\not$$\to\not{S}$, we define the Dirac operator
$D^{t}\mu=\sum_{i=1}^{7}e_{i}\cdot\nabla^{t}_{e_{i}}\mu,$ (5.6)
[…]